### v2

# LongReD: Mitigating Short-Text Degradation of Long-Context Large Language Models via Restoration Distillation

Abstract
Large language models (LLMs) have gained extended context windows through scaling positional encodings and lightweight continual pre-training. However, this often leads to degraded performance on short-text tasks, while the reasons for this degradation remain insufficiently explored. In this work, we identify two primary factors contributing to this issue: distribution drift in hidden states and attention scores, and catastrophic forgetting during continual pre-training. To address these challenges, we propose Long Context Pre-training with Restoration Distillation (LongReD), a novel approach designed to mitigate short-text performance degradation through minimizing the distribution discrepancy between the extended and original models. Besides training on long texts, LongReD distills the hidden state of selected layers from the original model on short texts. Additionally, LongReD also introduces a short-to-long distillation, aligning the output distribution on short texts with that on long texts by leveraging skipped positional indices. Experiments on common text benchmarks demonstrate that LongReD effectively preserves the model’s short-text performance while maintaining comparable or even better capacity to handle long texts than baselines. Our code is available at https://github.com/RUCAIBox/LongReD.
LongReD: Mitigating Short-Text Degradation of Long-Context Large Language Models via Restoration Distillation
Zican Dong1thanks: Equal Contribution., Junyi Li2∗, Jinhao Jiang1, Mingyu Xu3, Wayne Xin Zhao1thanks: Corresponding author.Bingning Wang3, Weipeng Chen3 1 Gaoling School of Artificial Intelligence, Renmin University of China 2 Department of Computer Science, National University of Singapore 3 Baichuan Inc. dongzican@ruc.edu.cn, junyi_cs@nus.edu.sg batmanfly@gmail.com, daniel@baichuan-inc.com
1 Introduction
Figure 1: Comparison of the original and long-context models via ABF Xiong et al. (2024) or PI Chen et al. (2023) on common short-text benchmarks.
Large language models (LLMs) have exhibited remarkable performance across a wide range of text and multimodal tasks Brown et al. (2020); OpenAI (2023); Zhao et al. (2023); Touvron et al. (2023); Dubey et al. (2024). However, their abilities to process long contexts are constrained by the positional encodings and the attention mechanism, which define the context window size based on the length of the pre-training data Su et al. (2024); Press et al. (2022). When the input text exceeds this context window, the model encounters out-of-distribution positional information, resulting in significant performance degradation Chen et al. (2023); Peng et al. (2023b); Dong et al. (2024).
To address the challenges of long context processing, various methods have been proposed to extend the context window of LLMs bloc97 (2023); Xiao et al. (2023); Xiong et al. (2024); Chen et al. (2023). A prominent approach involves scaling positional encodings combined with lightweight continual pre-training. Leveraging the properties of RoPE Su et al. (2024), these methods interpolate positional indices or adjust the RoPE base to prevent out-of-distribution rotary angles beyond the original context window Chen et al. (2023); Ding et al. (2024); Peng et al. (2023b); Xiong et al. (2024). Subsequently, lightweight continual pre-training adapts LLMs to the extended context and modified encodings, enabling context window up to 128K or even 1M tokens Xiong et al. (2024); Fu et al. (2024); Zeng et al. (2024). However, as shown in Figure 1, these long context window extension techniques come at the cost of degraded performance on short-text tasks Xiong et al. (2024). Yet, the root causes of this performance decline and potential mitigation strategies remain under-explored.
In this study, we aim to demystify the factors contributing to the performance degradation of short-text tasks after context window extension. Our analysis identifies two critical factors: distribution drift and catastrophic forgetting. The key findings of our work include: (1) Continual pre-training seeks to recover the original model’s internal distribution, but the restoration is inherently imperfect. (2) Distribution shift in hidden states potentially leads to performance degradation. (3) During continual pre-training, the performance on short-text tasks initially improves but subsequently declines as training progresses, highlighting the presence of catastrophic forgetting. (4) Replaying short text data is effective in mitigating forgetting and improving performance stability.
Based on our observations, we propose a novel approach called Long Context Pre-training with Restoration Distillation (LongReD) to mitigate the degradation in short-text capabilities of long-context LLMs. The central idea is that the short-text capacities of the extended model can be better preserved if it accurately simulates the original distributions before extension. To achieve this, in addition to typical long-text training, we introduce a short-text distillation objective, which employs the original model as a teacher to distill its hidden states on short texts into the extended model. This training objective minimizes distribution drift and alleviates catastrophic forgetting. Moreover, we propose a short-to-long distillation training objective to bridge the gap between short-text distillation and long-text training. In this setup, the original and extended models are fed with normal positional indices and skipped positional indices, respectively. By applying the distillation on the output distributions of the last layer, the short-text capacities can be effectively transferred and integrated into long-text processing.
To the best of our knowledge, this work represents the first systematic analysis of the reasons behind the performance degradation of long-context LLMs on short-text tasks. Furthermore, we propose a general method to mitigate this short-text degradation. To assess the effectiveness of our method, we extend the context window of Llama-3-8B and Mistral-7B-v0.3 and assess their performance on both short-text and long-text tasks. Experimental results demonstrate that our method preserves the original models’ performance on short-text tasks while maintaining or even improving their long-context modeling capabilities.
2 Background
2.1 Transformer
Owing to strong capabilities and scalability, Transformer decoders have served as the backbone of most LLMs Vaswani et al. (2017); Touvron et al. (2023); Dubey et al. (2024). Given a Transformer decoder with $L$ layers and $N$ heads and an input sequence $\bm{x}=\{x_{1},\dots,x_{T}\}$ consisting with $T$ tokens, the output of the $l$-th layer can be denoted as $\mathbf{H}_{l}=\{\bm{h}_{l,1},\dots,\bm{h}_{l,T}\}$. Each Transformer layer consists of a multi-head attention (MHA) module and a feed-forward network (FFN) module with residual connections connecting them, as shown by the following formula:
$\widetilde{\mathbf{H}}_{l}$ $=\mathrm{MHA}(\mathbf{H}_{l-1})+\mathbf{H}_{l-1},$ (1)
$\mathbf{H}_{l}$ $=\mathrm{FFN}(\widetilde{\mathbf{H}}_{l})+\widetilde{\mathbf{H}}_{l}.$ (2)
In the $i$-th head of the MHA module at the $l$-th layer, the hidden states are first projected into query, key, and value matrices, i.e., $\mathbf{Q}_{l}^{i}$, $\mathbf{K}_{l}^{i}$, and $\mathbf{V}_{l}^{i}$. Positional information is then incorporated into the query and key matrices through the RoPE with the rotation matrix $\mathbf{R}_{\theta}$ ($\theta$ is the RoPE base) Su et al. (2024). These matrices are subsequently processed through a dot product followed by a softmax operation to compute the attention scores $\mathbf{A}_{l}^{i}$. Finally, the value representations are weighted by the attention scores, and all attention heads are concatenated and projected to produce the attention output of the $l$-th layer as follows:
$\mathbf{A}_{l}^{i}$ $=\mathrm{Softmax}({\mathbf{Q}_{l}^{i}}^{\intercal}\mathbf{R}_{\theta}{\mathbf{K}_{l}^{i}}/\sqrt{d}),$ (3)
$\mathrm{MHA}(\mathbf{H}_{l-1})$ $=\mathrm{Concat}(\{\mathbf{A}_{l}^{i}\mathbf{V}_{l}^{i}\}_{i=1}^{N})\mathbf{W}^{O},$ (4)
where $d$ is the dimension of query and key, $\mathbf{W}^{O}$ is the projection matrix, and $\mathrm{Concat}$ is the concatenation of hidden states.
2.2 Measures of Distributional Discrepancy
After extending the context window, the parameters of LLMs undergo changes, resulting in a shift in the distribution of hidden states. To quantify such distributional discrepancy, we propose two measures, i.e., hidden state similarity and attention Kullback-Leibler (KL) divergence. Specifically, given the hidden states $\mathbf{H}_{l}$, $\mathbf{\widehat{H}}_{l}$ from the original and extended models, hidden state similarity measures the average cosine similarity between hidden states at the $l$-th layer across all positions. Similarly, attention KL divergence calculates the average KL divergences between the $i$-th attention head distributions $\mathbf{A}^{i}_{l}$, $\mathbf{\widehat{A}}^{i}_{l}$ from the original and extended models at the $l$-th layer. Higher hidden state similarity or lower attention KL divergence indicates less discrepancy between the two models. These metrics are formulated as follows:
$\operatorname{Sim}(\mathbf{H}_{l},\mathbf{\widehat{H}}_{l})$ $=\frac{1}{T}\sum_{t=1}^{T}\frac{{\bm{h}_{l,t}}^{\intercal}\bm{\widehat{h}}_{l,t}}{\lVert\bm{h}_{l,t}\rVert\lVert\bm{\widehat{h}}_{l,t}\rVert},$ (5)
$\operatorname{KL}(\mathbf{A}_{l}^{i},\widehat{\mathbf{A}}_{l}^{i})$ $=\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{t}a^{i}_{l,t,j}\log\frac{a^{i}_{l,t,j}}{\hat{a}^{i}_{l,t,j}}.$ (6)
3 Empirical Analysis
Prior work Xiong et al. (2024); Ding et al. (2024) demonstrated that the short-text capacities will degrade after lightweight continual pre-training for extending the context window of LLMs. However, the secrets behind the performance decline have not been fully explored. In this section, through conducting empirical analysis, we attribute the performance degradation to two factors, i.e., distribution drift and catastrophic forgetting.
3.1 Distribution Drift
Imperfect Distribution Restoration. In previous work (Chen et al., 2023; Ding et al., 2024), RoPE configurations are simply modified to extend the context window, undoubtedly leading to the change of distributions. Although continual pre-training is proposed to adapt LLMs to the extended context, a natural question is whether continual pre-training can eliminate the distributional discrepancy. To verify this, we modify RoPE base to $72e$ 7 and $81e$ 8, and train Llama-3-8B (with context window size of 8192 tokens) on 1B tokens with the sequence length of 32K and 128K, respectively, denoted as Llama-3-8B-32K and Llama-3-8B-128K. Besides, we select Llama-3-8B-Instruct-262K Pekelis et al. (2024), which is trained on Llama-3-8B-Instruct. To measure the distributional discrepancy, we employ these models before and after continual pre-training for inference on 1000 test samples with the length of 8192 tokens from SlimPajama Soboleva et al. (2023). The results of hidden state similarity and attention KL divergence are shown in Table 1. As can be seen, as the RoPE base increases, the extended LLMs before continual pre-training exhibit lower hidden state similarity and higher KL divergence from the original models. After training on long-context data, the similarity increases, and the KL divergence declines to some extent. However, despite the LLM striving to restore the inner distribution resembling the original model during continual pre-training, there still exists a certain degree of distributional discrepancy.
Model Base Simi $\uparrow$ KLD(×$\times 10^{-5}$) $\downarrow$
before after before after
Llama-3-8B-32K 2e7 0.92 0.95 3.51 1.51
Llama-3-8B-128K 1e8 0.83 0.94 6.74 1.68
Llama-3-8B-Ins-262K 2.8e8 0.76 0.91 8.75 2.50
Table 1: Results of hidden state similarity and KL divergence before and after continual pre-training.
Relationship between Distribution Drift and Performance Degradation. Owing to the distribution drift of the extended model, the inner working mechanisms may also change when processing the original short text data. To answer the question of whether the distributional discrepancy contributes to the performance decline in short text, we train the Llama-3-8B model using various context window extension methods and training strategies (detailed in Appendix D). Additionally, we incorporate existing open-source long-context models, which were extended from Llama-3-8B-Instruct by Gradient AI Pekelis et al. (2024). We compute the average hidden state similarity across all layers between these extended models and their original counterparts and assess their overall performance on the MMLU benchmark (mainly focused on short-text tasks) Hendrycks et al. (2021a). Then, we compute the MMLU performance preservation ratio of long-context models compared to the original models w.r.t. the hidden state similarity between them, as shown in Figure 2. Generally, extended models that maintain higher similarity to their original versions preserve more short-text performance. This finding highlights that distribution shifts in extended models are a significant factor contributing to performance degradation in short text.
Figure 2: Relationship between MMLU performance preservation of long-context models w.r.t. the hidden states similarity.
3.2 Catastrophic Forgetting
Catastrophic forgetting is always a critical problem during continual pre-training Wu et al. (2024b). Owing to the distributional difference of short texts and long texts, adapting for long context window may unavoidably leads to a trade-off on short texts. In this section, we examine the effect of the training steps and training data on the short-text performances to verify the forgetting phenomenons.
Effect of Training Steps. We train Llama-3-8B with the RoPE base of $72e$ 7 on the SlimPajama dataset and obtain checkpoints with different training steps, where the batch size is $64$ and the training length is 32K. Then, we evaluate the models on four short-text benchmarks, i.e., MMLU Hendrycks et al. (2021a), HumanEval Chen et al. (2021), TriviaQA Joshi et al. (2017), and PIQA  Bisk et al. (2020). Figure 3 presents the performance change with different training steps. It is clear that the performance is restored at the initial steps (usually less than 32 steps) in spite of fluctuations in some datasets. However, the performance gradually drops with the increase of training steps, which demonstrates that the continual adaptation for long context window will lead to catastrophic forgetting issues on short-text tasks.
Figure 3: Results of models with different training steps.
Effect of Training Data Mixture. Beyond the training steps, we also evaluate the effect of mixtures of training data. We replace half of the long text data with short text data with a length of 8K. We train Llama-3-8B on 1B tokens and evaluate them on different benchmarks. As shown in Table 2, the length of training data plays a critical role in downstream tasks. Introducing short-text data achieves better short-text performances than only training on long-text data. This verifies that long texts are critical for forgetting problems, which can be mitigated by replaying short texts.
Length MMLU HUMANEVAL PIQA TriviaQA
Long 62.0 14.02 74.10 70.67
Long+Short 62.5 16.46 78.24 72.82
Table 2: Results of models trained with data mixtures with different lengths.
4 Approach
Inspired by the above findings, we propose Long Context Pre-training with Restoration Distillation (LongReD), a novel context window extension framework to improve short-text capacities of long-context LLMs through decreasing distribution drift and mitigating catastrophic forgetting. Unlike only training on long texts, our strategy combines three different training objectives at each training step, i.e., long-text training, short-text distillation, and short-to-long distillation. For the three objectives, we employ three datasets $\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3}$ with different lengths $T_{l}$, $T_{s}(<T)$, and $T$, respectively, where $T$ and $T_{l}$ are the original and target long context window length, and $T_{s}$ is the short text length. The overall architecture is displayed in Figure 4.
Figure 4: Overview of our proposed Long Context Pre-training with Restoration Distillation (LongReD). The method consists of three parts, i.e., long-text training, short-text distillation, and short-to-long distillation.
4.1 Long Text Training
To extend the context window of an LLM, we follow previous work to first scale positional encodings via ABF Xiong et al. (2024) or PI Chen et al. (2023) and directly continually training on the long text dataset. Given any long text $\bm{x}$ from the dataset $\mathcal{D}_{1}$, the model $\Theta_{e}$ is trained with the language modeling objective to minimize the cross-entropy loss $\mathcal{L}_{long}$ as follows:
$\mathcal{L}_{long}=-\sum_{\bm{x}\in\mathcal{D}_{1}}\frac{1}{T_{l}}\sum_{t=1}^{T_{l}}\log\operatorname{Pr}(x_{t}|\bm{x}_{<t};\Theta_{e}).$ (7)
Through training on long texts, the model can learn to model long-term dependencies and adapt to the extended context window.
4.2 Short-Text Distillation
Section 3 demonstrates the negative effect of distribution drift and catastrophic forgetting on short-text capacities while only training on long texts. Though replaying short text data can partly mitigate the issue, the performance restoration is limited. Therefore, we propose short-text distillation, a knowledge distillation method to minimize the distribution discrepancy between the extended model $\Theta_{e}$ and the original model $\Theta_{o}$.
Distillation Layer Selection. To decrease the distribution discrepancy, we choose to distill the inner hidden states of LLMs. Yet, ensuring all the hidden states are the same as the original model is impossible and may significantly decrease the long text modeling capacities. Thus, we only select $M(<L)$ layers as the distill layers $\mathbf{L}_{M}=\{l_{1},\dots,l_{M}\},\forall l_{i}\in\{0,\dots,L\}$ and distill the outputs of these layers. Specifically, we propose an attention-based layer selection method where we first compute the KL divergence of attention scores between the extended and original models and then select layers with the largest KL divergences for distillation. We also select the critical last layer that directly determines the output Men et al. (2024).
Short-Text Distillation Loss. Given an input sample $\bm{x}$ from the short text dataset $\mathcal{D}_{2}$, we input the sample into the two models with its original positional indices $\bm{p}=[0,\dots,T_{s}-1]$. Then, we obtain the output hidden states of the selected layers of the two models, i.e., $\{\mathbf{H}_{l_{1};\bm{p},\Theta_{e}},\dots,\mathbf{H}_{l_{M};\bm{p},\Theta_{e}}\}$ and $\{\mathbf{H}_{l_{1};\bm{p},\Theta_{o}},\dots,\mathbf{H}_{l_{M};\bm{p},\Theta_{o}}\}$. We compute the cosine similarity between the hidden states of the two models at the same layer and position and sum up them as the short-text distillation loss $\mathcal{L}_{short}$, formulated as follows:
$\mathcal{L}_{short}=-\sum_{\bm{x}\in\mathcal{D}_{2}}\sum_{i=1}^{M}\operatorname{Sim}(\mathbf{H}_{l_{i};\bm{p},\Theta_{e}},\mathbf{H}_{l_{i};\bm{p},\Theta_{o}}).$ (8)
4.3 Short-to-Long Distillation
The two objectives of long-text training and short-text distillation separately optimize the model on long and short texts, resulting in a challenging performance trade-off between long and short texts. We argue that long and short texts share some inherent linguistic commonalities. Thus, we propose an additional short-to-long distillation to transfer the short-text capabilities to long texts, further bridging the gap between short and long texts.
Skipped Positional Indices. To distill short-text capabilities to long texts, we take the short text with the original context window size $T$ as input but employ the skipped positional indices method, which can simulate long-text positions and capture longer-term dependencies with short texts Zhu et al. (2024); Wu et al. (2024a). To be specific, for a sequence $\bm{x}$ drawn from $\mathcal{D}_{3}$, we first define a threshold $T_{b}$ to split the positional indices into three parts, i.e., $\bm{p}_{head}$, $\bm{p}_{mid}$, and $\bm{p}_{tail}$:
$\bm{p}$ $=\bm{p}_{head}\cup\bm{p}_{mid}\cup\bm{p}_{tail},$ (9)
$\bm{p}_{head}$ =$=\{0,\dots,T_{b}-1\},$ (10)
$\bm{p}_{mid}$ =$=\{T_{b},\dots,T-T_{b}-1\},$ (11)
$\bm{p}_{tail}$ =$=\{T-T_{b},\dots,T-1\}.$ (12)
Then, we modify the positional indices of the three segments. We keep $\bm{p}_{head}$ unchanged but modify the end positional index of $\bm{p}_{tail}$ to be equal to the target long text length $T_{l}$. For the middle segment, we employ uniform sampling or CREAM Wu et al. (2024a) which upsamples middle positions to determine the end positional index of $\bm{p}_{mid}$ as $T_{me}\in\{T-T_{b},\dots,T_{l}-T_{b}-1\}$. And the start positional index of $\bm{p}_{mid}$ is equal to $T_{me}$ minus the length of this middle segment $T-2T_{b}$, i.e., $T_{me}-T+2T_{b}$. The modified positional indices of the input text are formulated as follows:
$\hat{\bm{p}}$ $=\hat{\bm{p}}_{head}\cup\hat{\bm{p}}_{mid}\cup\hat{\bm{p}}_{tail},$ (13)
$\hat{\bm{p}}_{head}$ =$=\{0,\dots,T_{b}-1\},$ (14)
$\hat{\bm{p}}_{mid}$ =$=\{T_{me}-T+2T_{b},\dots,T_{me}\},$ (15)
$\hat{\bm{p}}_{tail}$ =$=\{T_{l}-T_{b},\dots,T_{l}-1\}.$ (16)
Details of the skipped positional indices methods are introduced in Appendix A.
Short-to-Long Distillation Loss. After obtaining skipped positional indices $\hat{\bm{p}}$, we input these indices into the extended model $\Theta_{e}$ while using the normal positional indices $\bm{p}$ for teacher model $\Theta_{o}$. Then, we obtain the outputs of the $L$-th last layer of both models and compute the cosine similarity between them. Then, we regard the negative cosine similarity as the loss to distill the output distribution into longer positions. The training objective $\mathcal{L}_{s2l}$ can be represented as follows:
$\mathcal{L}_{s2l}=-\sum_{\bm{x}\in\mathcal{D}_{3}}\operatorname{Sim}(\mathbf{H}_{L;\hat{\bm{p}},\Theta_{e}},\mathbf{H}_{L;\bm{p},\Theta_{o}}).$ (17)
Notably, we only distill the output of the last layer instead of selected layers in short-text distillation to avoid the disturbance of latent positional information in the middle layers.
4.4 Joint Training Objective
Finally, we aggregate losses from these three training objectives. For each batch of training data, we sample data points from the three datasets with a fixed ratio. Then, we independently compute loss through the three different objectives on the corresponding data and sum them up to balance the long and short text capacities. The final loss $\mathcal{L}_{final}$ can be represented as follows:
$\mathcal{L}_{final}=\mathcal{L}_{long}+\alpha_{1}\mathcal{L}_{short}+\alpha_{2}\mathcal{L}_{s2l},$ (18)
where $\alpha_{1}$ and $\alpha_{2}$ are hyper-parameters that control the proportion of the three parts.
5 Experiment
Model CW PE Data Method General Coding Math RC Common Short Avg. RULER Avg.
Llama-3 -8B 8K - - - 68.18 41.20 30.12 63.78 72.54 55.16 - -
32K ABF Long CPT 65.39 29.32 29.50 60.98 69.83 51.00 82.80 56.30
32K ABF Mix CPT 66.04 27.40 29.32 34.65 71.88 45.86 85.04 52.39
32K ABF Mix LongReD-C 67.65 39.90 31.52 62.84 72.35 54.85 84.98 59.87
32K ABF Mix LongReD-U 67.05 39.71 30.90 62.51 72.61 54.56 84.08 59.48
32K PI Long CPT 63.84 27.13 26.18 61.95 67.71 49.36 82.17 54.83
32K PI Mix CPT 64.22 27.29 29.16 36.63 71.09 45.68 81.63 51.67
32K PI Mix LongReD-C 66.45 38.39 30.78 62.79 71.30 53.94 81.19 58.48
32K PI Mix LongReD-U 66.34 38.83 31.68 62.76 72.36 54.39 79.30 58.54
128K ABF Long CPT 64.06 30.04 27.72 60.49 68.37 50.14 69.70 53.40
128K ABF Mix CPT 66.39 14.03 29.02 35.28 71.91 43.33 69.64 47.71
128K ABF Mix LongReD-C 65.99 40.03 30.26 62.39 71.50 54.03 64.93 55.85
128K ABF Mix LongReD-U 65.61 39.87 29.67 62.47 71.64 53.85 68.41 56.28
Mistral -7B-v0.3 32K - - - 65.36 33.20 26.98 68.07 61.86 51.09 - -
128K ABF Long CPT 55.29 26.37 15.79 52.93 53.07 40.68 44.63 41.35
128K ABF Mix CPT 55.99 26.36 15.30 60.21 52.74 40.66 45.76 41.51
128K ABF Mix LongReD-C 61.45 26.66 19.88 65.25 58.26 46.30 58.37 48.31
128K ABF Mix LongReD-U 62.42 29.64 19.56 66.86 59.98 47.69 53.60 48.68
Table 3: Comparison of performances of short-text and long-text benchmarks of our methods with other baselines. CW denotes the context window length, PE denotes the scaling method of RoPE, RC denotes reading comprehension, Common denotes commonsense question answering, Short Avg. denotes averaged scores on short benchmarks, and Avg. denotes averaged scores of all benchmarks. LongReD-C and LongReD-U denote our method with different skipped positional indices methods, i.e., CREAM and Uniform Sampling.
5.1 Experimental Settings
Pre-training Setup. In the experiments, we choose Llama-3-8B Dubey et al. (2024) and Mistral-7B-v0.3 Jiang et al. (2023) to evaluate the effectiveness of our methods on context window extension. We sample the long text dataset $\mathcal{D}_{1}$ and short-to-long dataset $\mathcal{D}_{3}$ from SlimPajama following the setting of Fu et al. (2024). Inspired by Gao et al. (2024), we employ higher-quality data as the short text dataset $\mathcal{D}_{2}$ with the length of 1K. Then, we trained our model on these three datasets with a token quantity ratio of 4:3:1 and a total of 1B tokens. The number of distillation layers $M$ is set to 3 for extending Llama-3-8B to 128K context window size and 6 for other settings for better balancing the short and long text modeling performances. The details of training and parameter configurations and data mixture are displayed in Appendix B.
Evaluation Benchmarks. To thoroughly evaluate the short-text performance of long-context models, we choose 17 benchmarks covering 5 capacities, i.e., general, coding, math, reading comprehension, and commonsense question answering. We also select RULER Hsieh et al. (2024) to assess the long text processing abilities. The evaluation datasets and details are shown in Appendix C.
Baselines. In our experiments, we choose two baseline methods for comparative analysis: (1) Long+CPT: Only continual pre-training on long text datasets $\mathcal{D}_{1}$; (2) Mix+CPT: Continually pre-training on a mixed-length dataset ($\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\mathcal{D}_{3}$), keeping the training data the same as our method.
5.2 Main Results
Table 3 presents the performance comparison between our proposed methods and various baselines on both short-text and long-text tasks. Detailed results for each dataset can be found in Appendix F.
First, LongReD performs better on short-text tasks and achieves competitive long-text modeling performances compared to baseline methods. Across various scaling methods and target context window sizes, LongReD consistently outperforms most baselines on short-text benchmarks. Specifically, when extending Llama-3-8B to a 32K-token context window using ABF and LongReD-C, the short-text capabilities are retained up to $99.4\%$ of the original model’s performance, compared to $92.5\%$ achieved by naive long-text training. Furthermore, our methods demonstrate competitive performance with Llama-3-8B and superior performance with Mistral-7B-v0.3 on the RULER benchmark for long-text processing, when compared to continual pre-training approaches. These compelling results robustly validate the efficacy of our proposed methodology.
Second, the skipped positional indices method is crucial for long-text performances. Comparing LongReD with CREAM and uniform sampling methods, we find that the short-text performances are similar while the long-text performances on RULER are largely different. When the extension ratio is relatively small (e.g.,, 4 times), CREAM demonstrates superior performance by effectively mitigating the lost-in-the-middle problem. However, on the length of 128K tokens, the model trained with CREAM performs largely worse than that with uniform sampling. We hypothesize that the excessive focus on the middle positions introduced by CREAM results in some positions not being adequately trained.
Finally, our method benefits from the scaling techniques of positional encodings. When scaling LLaMA-3-8B to 32K with either ABF or PI, our LongReD consistently achieves superior short-text performances. However, a comparison between models with different extension methods reveals that PI exhibits inferior performance to ABF on both short-text and long-text tasks when combined with LongReD. We hypothesize that this performance gap arises due to the significant distribution discrepancy introduced by PI, which poses greater challenges for our method to mitigate effectively.
5.3 Detailed Analysis
In this section, we conduct further detailed analysis to investigate our approach.
Ablation Study. In addition to long-text training, our method incorporates both short-text distillation and short-to-long distillation. To evaluate the individual contributions of these two components, we employ half of the training data for long-text training and another half data for short-text distillation or short-to-long distillation. Moreover, we also evaluate the effect of the hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on both short and long text tasks. The results are compared against the LongReD-C method with $\alpha_{1}=5,\alpha_{2}=10$, as summarized in Table 4. The results indicate that excluding short-text distillation significantly degrades the model’s performance on short-text tasks. Conversely, omitting the short-to-long distillation stage results in noticeable declines in long-text performance. In addition, The adaptation of the hyper-parameters is essential. Increasing or decreasing $\alpha_{2}$ will lead to performance degradation on long-text tasks, while slightly decreasing $\alpha_{1}$ will lead to better long-text performances. When either $\alpha_{1}$ or $\alpha_{2}$ is set to an extremely large value, the model’s performance on long-context tasks drops precipitously.
$\alpha_{1}$ $\alpha_{2}$ General Code Math Common RC RULER
5 10 67.65 39.90 31.52 72.35 62.84 84.98
- 15 66.54 39.93 31.68 70.64 62.47 85.48
5 - 67.39 41.25 30.78 71.50 62.25 83.61
5 100 66.96 40.47 33.12 71.86 62.47 80.40
5 30 67.00 40.14 31.59 72.26 62.29 83.82
5 15 67.20 38.20 31.46 72.39 62.55 84.87
5 5 66.91 41.84 32.67 72.36 62.42 83.75
100 10 65.33 37.81 29.24 72.00 61.85 80.19
2 10 67.02 41.10 31.83 72.25 63.04 86.03
Table 4: Ablation results on five tasks. “-” denotes the objective is not employed.
Distillation Layers. We explore the impact of distillation layers by varying the number of distilled layers and their selection strategies. In comparison to our method of selecting six layers via KL divergence (denoted as KL(6)), we design three variants: (1) All: distilling all layers; (2) Last: distilling only the last layer; and (3) Uniform(6): uniformly selecting six layers. The results are summarized in Table 5. We can observe that distilling either all the layers or the last layer leads to worse long-text performance and general capacities than selecting six layers based on KL divergence. Furthermore, uniformly selected layers consistently exhibit inferior performance to those selected via attention KL divergence across the long-text tasks for the same number of distilled layers. These findings underscore the critical role of layer selection methods and distillation granularity in the performances of long-context LLMs.
Layers General Code Math Common RC RULER
KL(6) 67.65 39.90 31.52 72.35 62.84 84.98
Uniform(6) 66.86 40.24 32.19 72.33 60.27 82.53
All 66.89 39.63 31.80 72.33 62.61 81.47
Last 66.40 40.49 30.69 72.11 63.19 82.24
Table 5: Results with different distillation layers.
Distillation Length. In our experiments, we set the length of short-text distillation $T_{s}$ as $1024$. To explore the effect of distillation length, we set $T_{s}$ as $2048$ and $8192$, and check the performance changes. As shown in Table 6, increasing the distillation length will harm the model’s ability to handle long texts while the short-text capacities are generally the same. Overfitting to the long distillation length will destroy the implicit positional information in hidden states, which may be the main reason for the performance drop (details are shown in Appendix E).
$T_{s}$ General Code Math Common RC RULER
1024 67.65 39.90 31.52 72.35 62.84 84.98
2048 66.80 40.13 31.01 72.63 62.32 83.20
8192 67.43 39.26 31.41 71.75 63.04 75.54
Table 6: Results with varying distillation length $T_{s}$.
Comparison with Continual Learning Methods. Beyond simple continual pre-training, we also compare LongReD with two continual learning methods that can effectively mitigate catastrophic forgetting issues: (1) Model Merging: Following previous work Hu et al. (2024b), we average the checkpoints of the original model and the extended model; (2) Parameter-Efficient Tuning: Since modifying RoPE directly affects attention computation without altering the knowledge stored in MLP parameters, we only tune the parameters in the attention blocks. We train Llama-3-8B on a 32K context length with these methods and evaluate their performances. As shown in Table 7, though these methods can achieve better performances than naive continual pre-training, our method consistently performs better on most short-context capacities evaluations.
Method General Code Math Common RC RULER
LongReD-C 67.51 39.90 31.52 72.35 62.84 84.98
Merging 65.92 35.04 32.38 71.61 62.65 83.38
PET 66.82 39.20 30.16 70.78 62.26 85.86
CPT 65.39 29.32 29.50 69.83 60.98 82.80
Table 7: Comparison with continual learning methods, where Merging denotes model merging and PET denotes parameter-efficient tuning.
6 Related Work
Context Window Extension. LLMs typically have constrained context windows owing to positional encodings and attention mechanisms Press et al. (2022); Han et al. (2023). In order to meet the growing demand for processing long texts, various methods have been proposed to extend the context windows of LLMs based on modification of RoPE Su et al. (2024). Positional Interpolation down-scale positional indices to prevent out-of-distribution positions. In addition, NTK bloc97 (2023); emozilla (2023), ABF Xiong et al. (2024), Yarn Peng et al. (2023b), and LongRoPE Ding et al. (2024) modify the base of RoPE to control the rotation angles of different dimensions. After modifying positional encodings, LLMs are lightweightly continued pre-training on long texts Fu et al. (2024). Through this pipeline, the context windows of LLMs can be extended to 128K or even 1M tokens Dubey et al. (2024); Zeng et al. (2024). Typically, they can achieve utilization of long texts within their context windows at the cost of degradation of general capacities on short texts Dubey et al. (2024). Our work mainly focuses on the reasons for the degradation of short-text capacities and approaches to mitigating the decline.
Knowledge Distillation. Knowledge distillation serves as an approach to transferring knowledge and abilities of teacher models to student models Hinton et al. (2015); Xu et al. (2024). During the era of LLMs, leveraging black-box and powerful LLMs, e.g., GPT-4 OpenAI (2023), to generate instructions and even fine-grained reasoning path to training smaller LLMs is a common and effective method to enhance general and domain-specific capacities Chiang et al. (2023); Peng et al. (2023a); Mukherjee et al. (2023); Ho et al. (2023). In addition, distilling the inner distribution of white-box LLMs can be another effective way. Divergence-based distillation methods minimize the divergence of the output probability distribution of teacher and student models Gu et al. (2024); Agarwal et al. (2024); Jiang et al. (2024). Similarity-based distillation methods align the intermediate hidden states, enabling the student models to process input like the teacher models Liang et al. (2023); Muralidharan et al. (2024). Different from these methods of transferring knowledge of powerful models to small or compressed models, our method aims to restore and preserve the original capacities of LLMs during the context window extension stage.
7 Conclusion
In this paper, we analyzed two reasons for the performance degradation in short-text tasks after context window extension, i.e.,distribution drift and catastrophic forgetting. Based on the observations, we proposed an effective training method to better preserve short-text capacities, named Long Context Pre-training with Restoration Distillation (LongReD). Besides continual pre-training on long texts, LongReD introduced two additional training objectives: short-text distillation and short-to-long distillation. Experiments demonstrated that our method could achieve better performances on common short-text benchmarks while achieving comparable long-text modeling abilities on the RULER benchmark compared with continual pre-training. In future work, we will explore how to directly integrate continual training with the distillation of the original model on long texts, rather than applying them separately to texts of different lengths.
Limitations
In this paper, we present a novel perspective on the extension of context windows and the associated decline in the general capabilities of LLMs. We also propose a method to preserve general capabilities while enhancing long-context abilities. However, a notable degradation in short-text performance after lightweight continual pretraining on only several billion tokens. Xiong et al. (2024) and Dubey et al. (2024) have shown that, with training on more than 100 billion tokens, some capabilities on short texts of models may remain largely unaffected or even improve. Furthermore, our proposed method offers a general training strategy that is compatible with different positional encoding extension techniques. A more refined extension method that minimizes disruption to the model’s distribution could further enhance the effectiveness of our approach, a topic we leave for future investigation.
Appendix A Skipped Positional Indices
Skipped positional indices are widely employed for effective long-context training. The method simulates long-distance dependency via modifying the positional indices. In this section, we introduce two skipped positional indices employed in our methods.
A.1 CREAM
Continuity-Relativity indExing with gAussian Middle (CREAM) Wu et al. (2024a) is a method that modifies position indices to simulate long positions. The main idea is to make the model better focused on the middle part of the long sequences.
As described in Section 4.3, the positional indices are first split into three non-overlapped parts, i.e., $\bm{p}_{head}$, $\bm{p}_{tail}$, and $\bm{p}_{tail}$, where the length of the head and the tail are $T_{b}$. There are two segmentation methods and we randomly select one method to determine the number of $T_{b}$. To achieve the continuity of the middle part, $T_{b}$ is set as a small number, where we set it to four times the context window extension factor. To learn more relative positions, $T_{b}$ is set as one-third of the input length $T/3$.
To better concentrate on the middle positions, the method proposes a truncated Gaussian function to determine the end position of the middle part $p_{e}$. Initially, the method defines a Gaussian distribution $f(x)$ and its cumulative distribution function $F(x)$:
$f(x)$ $=\frac{1}{\sigma\sqrt{2}\pi}\exp(-\frac{(x-\mu)^{2}}{2\sigma^{2}}),$ (19)
[…]
where $\mu=1+T_{l}/T$ and $\sigma=3$. Then, the method samples 1000 points $x_{i}$ uniformly from 1 to extension factor $T_{l}/T$ and calculates their cumulative distribution function $F(x_{i})$. After that, the method samples $F(u)$ from the Gaussian distribution and uses interpolation and inverse transformation to obtain the corresponding $u$. The number $u$ then rounds to the nearest integer $\alpha$.
$u$ $=x_{i-1}+\frac{(x_{i}-x_{i-1})(F(u)-F(x_{i-1}))}{F(x_{i})-F(x_{i-1})},$ (21)
$\alpha$ $=\operatorname{round}(u),$ (22)
where $F(x_{i})$ and $F(x_{i-1})$ are the two nearest numbers to $F(u)$ and $F(u)$ is sampled uniformly from 0 to 1. Finally, we sample an integer as the end of the middle part:
$T_{me}$ $\sim\operatorname{Uniform}(T_{b}+\alpha(T_{l}-2T_{b}),\alpha T-T_{b}-1),$ (23)
$T_{mb}$ $=T_{me}+2T_{b}-T.$ (24)
A.2 Uniform Sampling
In the uniform sampling method, we keep the same as the CREAM method except the sampling method of the positions of the middle part. Instead, we employ a uniform sampling method to determine the end positions of the middle part, as formulated as follows:
$T_{me}$ $\sim\operatorname{Uniform}(T-T_{b},T_{l}-T_{b}-1),$ (25)
$T_{mb}$ $=T_{me}+2T_{b}-T,$ (26)
where $\operatorname{Uniform}$ denotes uniformly sampling from this range.
Appendix B Training Details
B.1 Training Datasets
In our proposed method, we employ three datasets: $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$, which are specifically designed for long-text training, short-text distillation, and short-to-long distillation, respectively. Following the sampling strategy of Fu et al. (2024), we sample data in the same proportion as Llama-2 from the SlimPajama datasets Soboleva et al. (2023), tokenize the samples, and pack them to the target length. To enhance the preservation of short-text capabilities during the short-text distillation phase, we adopt the approach proposed by Gao et al. (2024), utilizing higher-quality datasets. Specifically, we select Fineweb-Edu Penedo et al. (2024), Stack-v2 Lozhkov et al. (2024), the mathematical subset of Fineweb-Edu Penedo et al. (2024), Proof-Pile-2 Azerbayev et al. (2024), PG19 Rae et al. (2020), Arxiv Soboleva et al. (2023), and Wikipedia Soboleva et al. (2023). These datasets are carefully chosen to balance quality and diversity. Additionally, we incorporate instruction-tuned datasets that have been demonstrated to be beneficial for model performance Hu et al. (2024a), including Tulu-v2-sft-mixture [^1] https://huggingface.co/datasets/allenai/tulu-v2-sft-mixture, MathInstruct [^2] https://huggingface.co/datasets/TIGER-Lab/MathInstruct, WizardLM-evol-instruct-V2333https://huggingface.co/datasets/WizardLMTeam/WizardLM_evol_instruct_V2_196k, and Magicoder-Evol-Instruct-110K [^4] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K. The detailed proportions of these datasets utilized in our method are summarized in Table 8.
Dataset Data Mixture
$\mathcal{D}_{1}$
Wikipedia: 0.034, CommonCrawl: 0.534,
StackExchange: 0.032, C4: 0.266,
Github: 0.050, ArXiv: 0.043, Book: 0.041
$\mathcal{D}_{3}$
Wikipedia: 0.034, CommonCrawl: 0.534,
StackExchange: 0.032, C4: 0.266,
Github: 0.050, ArXiv: 0.043, Book: 0.041
$\mathcal{D}_{2}$
Fineweb-Edu: 0.476, Fineweb-Edu-Math: 0.722
Proof-Pile-2: 0.230, Stack-v2: 0.095, PG19: 0.095,
Wikipedia: 0.096, ArXiv: 0.095, Instructions: 0.047
Table 8: Data Mixture of three datasets.
B.2 Evaluated Models
We select Llama-3-8B and Mistral-7B-v0.3 as our evaluated models. For Llama-3-8B, we extend its context window to 32K and 128K, with the ABF and PI methods. For Mistral-7B-v0.3, we extend its context window to 128K with ABF method. The configuration and training length of these models are displayed in Table 9.
Model CW PE Ratio $L_{\mathcal{D}_{1}}$ $L_{\mathcal{D}_{2}}$ $L_{\mathcal{D}_{3}}$
Llama-3-8B 32K ABF 2e7 32000 1024 8192
Llama-3-8B 32K PI 4 32000 1024 8192
Llama-3-8B 128K ABF 1e8 128000 1024 8192
Mistral-7B-v0.3 128K ABF 2e7 128000 1024 32768
Table 9: Configuration of models and length of datasets. CW denotes the context window of target models, PE denotes the scaling methods of RoPE, Ratio denotes the RoPE theta for ABF methods, and the interpolation ratio for PI methods. $L_{\mathcal{D}_{1}}$, $L_{\mathcal{D}_{2}}$, and $L_{\mathcal{D}_{3}}$ denotes the length of the three datasets.
B.3 Training Configurations
We employ ring flash attention Liu et al. (2023) in EasyContext framework Zhang et al. (2024) (Apache-2.0 License) to train our models with 8 A800 GPUs. The learning rate is fixed as 2e-5 without warmup and the training tokens of each batch are about 2M tokens. We employ AdamW optimizer Loshchilov and Hutter (2019) with the weight decay of 0.1, $\beta_{1}$ of 0.9, and $\beta_{2}$ of 0.95. The hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ are set as $\alpha_{1}=5,\alpha_{2}=10$ for 32K, and $\alpha_{1}=2,\alpha_{2}=15$ for 128K. All the models are trained with 512 steps with about 1B tokens.
B.4 Training Costs
We report the training costs of our method and other baselines. All the models are trained on 8 A800 GPUs with ring attention. The time costs for training these models are shown in Table 10. Compared with baselines, our methods will cost about $10\%$ extra computation than the only continually pre-training on the same data. Additionally, comparing only training on long texts, LongReD costs similar time in 32K while saving about $20\%$ computations in 128K settings.
Model CW Data Method Time(h)
Llama-3-8B 32K Long CPT 22.4
32K Mix CPT 19.9
32K Mix LongReD 22.4
128K Long CPT 42.8
128K Mix CPT 30.4
128K Mix LongReD 33.4
Table 10: Training time of Llama-3-8B with different methods and target context windows.
Appendix C Evaluation Details
To evaluate the performance of LLMs on short-text and long-text tasks, we employ multiple benchmarks. In this section, we introduce the datasets for evaluating different capacities:
  • General: MMLU Hendrycks et al. (2021a), BBH Suzgun et al. (2023), LAMBADA Paperno et al. (2016)
  • Math: MATH Hendrycks et al. (2021b), GSM8K Cobbe et al. (2021)
  • Coding: HumanEval Chen et al. (2021), MBPP Austin et al. (2021)
  • Reading comprehension: Squadv2 Rajpurkar et al. (2016), Quac Choi et al. (2018), TriviaQA Joshi et al. (2017), Drop Dua et al. (2019), BoolQ Clark et al. (2019)
  • Commonsense question answering: Openbookqa Mihaylov et al. (2018), Commonsenseqa Talmor et al. (2019), ARC-C Clark et al. (2018), SIQA Sap et al. (2019), PIQA Bisk et al. (2020)),
  • Long text processing: RULER Hsieh et al. (2024)
we also present details of our evaluation configurations, as shown in Table 11.
Dataset CoT #Shots Metric
MMLU $\times$ 5 Probability
BBH 3 EM
Lambada $\times$ 0 accuracy
MBPP $\times$ 3 pass@1
HumanEval $\times$ 0 pass@1
GSM8K 4 accuracy
MATH 4 accuracy
PIQA $\times$ 0 Probability
SIQA $\times$ 0 Probability
OpenBookQA $\times$ 0 Probability
ARC(C) $\times$ 25 Probability
CommonSenseQA $\times$ 7 PPL
Quac $\times$ 1 F1
TriviaQA $\times$ 5 F1
DROP $\times$ 3 F1
BoolQ $\times$ 3 F1
Squad-v2 $\times$ 1 F1
RULER $\times$ 0 Accuracy
Table 11: Configurations of evaluated benchmarks. CoT denotes employing the Chain-of-Thought prompt, #Shot denotes the number of shots in prompts, and Metric denotes the evaluation metric for the benchmark.
Model RoPE CW Base Model Tokens MMLU MMLU Ratio Simi
GradientAI Models
Llama-3-8B-Ins-G-262k ABF(2.8e8) 262K - - 61.9 0.945 0.920
Llama-3-8B-Ins-G-1048k ABF(3.6e9) 1048K - - 59.7 0.911 0.900
Llama-3-8B-Ins-G-4194k ABF(4.5e10) 4194K - - 56.1 0.856 0.872
Our Models (Directly Extension)
Llama-3-32K-5e6 ABF(5e6) 32K Llama-3-8B 256M 62.2 0.960 0.943
Llama-3-32K-5e7 ABF(5e7) 32K Llama-3-8B 256M 62.0 0.957 0.934
Llama-3-32K-5e8 ABF(5e8) 32K Llama-3-8B 256M 60.9 0.940 0.924
Llama-3-32K-5e9 ABF(5e9) 32K Llama-3-8B 256M 60.4 0.932 0.915
Llama-3-32K-5e10 ABF(5e10) 32K Llama-3-8B 256M 59.8 0.923 0.908
Llama-3-32K-5e20 ABF(5e20) 32K Llama-3-8B 256M 56.5 0.872 0.856
Llama-3-32K-PI4 PI(4) 32K Llama-3-8B 256M 60.9 0.940 0.917
Llama-3-32K-PI16 PI(16) 32K Llama-3-8B 256M 56.1 0.866 0.862
Our Models (Gradually Extension)
Llama-3-16K-5e6 ABF(5e6) 16K Llama-3-8B 256M 62.3 0.961 0.946
Llama-3-32K-5e7 ABF(5e7) 32K Llama-3-16K-5e6 256M 60.7 0.938 0.927
Llama-3-64K-5e8 ABF(5e8) 64K Llama-3-32K-5e7 256M 60.3 0.930 0.912
Llama-3-128K-5e9 ABF(5e9) 128K Llama-3-64K-5e8 256M 57.2 0.883 0.901
Table 12: Details of evaluated models for the analysis of the relationship between distribution drift and short-text performances.
The details of metrics employed to evaluate the performance are described as follows:
  • Probability: For a choice task, all the choices are provided in the prompt and the Probability of each choice is calculated. The choice with the maximum Probability will be compared with the correct choice and the accuracy is reported.
  • PPL: For a choice task, the choices are not provided. Instead, the content of each choice is input after the question, and the average perplexity of each choice is calculated. The choice with the lowest PPL is selected as the correct answer.
  • Pass@1: For coding tasks, the model is asked to give one solution. Then, we evaluate the solution with the unit testing and count the numbers that pass the testing.
  • EM: For the question-answering task, the exact matching metric measures whether the output is the same as the golden answer.
  • F1: For the question-answering task, the F1 first splits the golden answer and output of the model into words, and calculates the F1 score of these word lists.
  • Accuracy: Accuracy measures the number of correctly answered samples out of all the samples. For RULER, the sample is correct if the answer is in the output or each part of the answer (only for qa1 and qa2).
Appendix D Models for Evaluating Distribution Drift
To evaluate the relationship between the distribution drift and the short performance, we first employ open-sourced three long-context models from Gradient AI Pekelis et al. (2024), which are continually pre-trained from Llama-3-8B-Instruct with only several hundred million tokens. Then, we train Llama-3-8B with ABF and PI methods with different extension ratios on 256M tokens with a length of 32K. We also follow Pekelis et al. (2024), gradually extending the context window length to twice its original length. For each extension, we train the model with 256M tokens and enlarge the RoPE theta ten times. The full model list and their MMLU scores and hidden states similarity can be seen in Table 12. Typically, a larger extension ratio will cause a greater shift in the model’s distribution, further leading to a decline in the model’s performance on short texts. We provide theoretical proof of the impact of the RoPE bases on the drifts of attention distributions.
D.1 Proof of RoPE Base Impact on Attention Scores
Given a query $\mathbf{q}^{m}$ and a key $\mathbf{k}^{n}$, their inner product with RoPE encoding can be expressed as:
$(\mathbf{R}^{d}_{\Theta,m}\mathbf{q}^{m})^{\top}(\mathbf{R}^{d}_{\Theta,n}\mathbf{k}^{n})$ (27)
$=$ $\mathrm{Re}\left[\sum_{i=0}^{d/2-1}\mathbf{q}^{m}_{[2i:2i+1]}{\mathbf{k}^{n}_{[2i:2i+1]}}^{*}e^{\mathrm{i}(m-n)\theta_{i}}\right]$ (28)
where $[2i:2i+1]$ denotes the $i$-th subspace. Following the definition from Su et al. (2024), we let $h_{i}=\mathbf{q}^{m}_{[2i:2i+1]}{\mathbf{k}^{n}_{[2i:2i+1]}}^{*}$ and $S^{\Theta}_{j}=\sum_{k=0}^{j-1}e^{\mathrm{i}(m-n)\theta_{k}}$. We extend these definitions with $h_{d/2}=0$ and $S^{\Theta}_{0}=0$. Applying Abel transformation, we rewrite the summation:
$\sum_{i=0}^{d/2-1}h_{i}e^{\mathrm{i}(m-n)\theta_{i}}$ $=\sum_{i=0}^{d/2-1}h_{i}(S^{\Theta}_{i+1}-S^{\Theta}_{i})$ (29)
$=-\sum_{i=0}^{d/2-1}S^{\Theta}_{i+1}(h_{i+1}-h_{i})$ (30)
Theorem. Let $\Theta=b^{-2i/d}$ and $\Theta^{\prime}={b^{\prime}}^{-2i/d}$ be two RoPE bases with $b>b^{\prime}>1$. For fixed $\mathbf{q}^{m}$ and $\mathbf{k}^{n}$, their attention score difference satisfies:
$\left|\textit{Attn}(b)-\textit{Attn}(b^{\prime})\right|\leq C\cdot\sum_{i=0}^{d/2-1}\left(|S^{\Theta}_{i+1}|+|S^{\Theta^{\prime}}_{i+1}|\right)$
where $C=\max\limits_{0\leq i<d/2}|h_{i+1}-h_{i}|$.
Proof. The attention scores under different bases are:
$\textit{Attn}(b)=\mathrm{Re}\left[-\sum_{i=0}^{d/2-1}S^{\Theta}_{i+1}(h_{i+1}-h_{i})\right],$ (31)
$\textit{Attn}(b^{\prime})=\mathrm{Re}\left[-\sum_{i=0}^{d/2-1}S^{\Theta^{\prime}}_{i+1}(h_{i+1}-h_{i})\right]$ (32)
Their difference satisfies:
$\left|\textit{Attn}(b)-\textit{Attn}(b^{\prime})\right|$ (33)
$\leq$ $\left|\sum_{i=0}^{d/2-1}\left(S^{\Theta}_{i+1}-S^{\Theta^{\prime}}_{i+1}\right)(h_{i+1}-h_{i})\right|$ (34)
$\leq$ $\sum_{i=0}^{d/2-1}\left|S^{\Theta}_{i+1}-S^{\Theta^{\prime}}_{i+1}\right|\cdot\left|h_{i+1}-h_{i}\right|$ (35)
$\leq$ $C\sum_{i=0}^{d/2-1}\left(|S^{\Theta}_{i+1}|+|S^{\Theta^{\prime}}_{i+1}|\right)$ (36)
where the last inequality follows from $|S^{\Theta}_{i+1}-S^{\Theta^{\prime}}_{i+1}|\leq|S^{\Theta}_{i+1}|+|S^{\Theta^{\prime}}_{i+1}|$ by triangle inequality.
To analyze the expected difference across positions, consider:
$\mathbb{E}\left[\left|\textit{Attn}(b)-\textit{Attn}(b^{\prime})\right|\right]$ (37)
$\leq$ $\frac{1}{N}\sum_{q,k,t}C^{q,k}\sum_{i=0}^{d/2-1}\left(|S^{\Theta,t}_{i+1}|+|S^{\Theta^{\prime},t}_{i+1}|\right)$ (38)
$\leq$ $\frac{C}{N}\sum_{t=0}^{T}(T-t)\sum_{i=0}^{d/2-1}\left(|S^{\Theta,t}_{i+1}|+|S^{\Theta^{\prime},t}_{i+1}|\right)$ (39)
Base 5e5 5e6 5e7 5e8 5e9 5e10 5e11 5e12 5e13
$B(\Theta)$ 1.3e6 1.5e6 1.9e6 2.3e6 3.0e6 3.3e6 3.6e6 3.9e6 4.1e6
Table 13: Upper bound of different RoPE bases.
where $T$ is the maximum sequence length and $t=|m-n|$ is the relative position. Let $B(\Theta)=\sum_{t=0}^{T}(T-t)\sum_{i=0}^{d/2-1}|S^{\Theta,t}_{i+1}|$. Previous work has demonstrated that RoPE has long-term decay and a large base will lead to slow attention decay. For bases $b_{3}>b_{2}>b_{1}$, we have $B(\Theta_{3})>B(\Theta_{2})>B(\Theta_{1})$ due to slower attention decay with larger bases. As shown in Table 13, empirical measurements on Llama-3-8B confirm that $B(\Theta)$ increases with $b$, proving that larger bases induce greater attention score changes during context window extension.
Appendix E Analysis of Distillation Length
As discussed in Section 5.3, increasing the distillation length will harm the long text modeling capacities. To explore the reason for the performance decline, we employ positional vectors Dong et al. (2024), a method to extract latent positional information from the hidden states. Specifically, given the hidden states from different samples, we average the hidden states at the same position to obtain vectors that representing positional information, denoted as positional vectors:
$\bm{p}_{l,i}=\frac{1}{N}\sum_{s=1}^{N}\bm{h}_{l,i}^{s}$ (40)
Then, we compute the cosine similarities of positional vectors with different places. All the positional vectors are calculated on samples from SlimPajama Soboleva et al. (2023) with a length of 32000 tokens. The similarity matrices of the positional vectors of models with distillation lengths of 1024, 2048, and 8192 as well as the models trained with only long texts are shown in Figure 5. We can observe that a large distillation length will cause a significant discontinuity in the implicit positional information inside and outside the original context window.
Figure 5: Simialrity matrices of positional vectors inside and outside the context window.
Appendix F Results Details
In this section, we display the details of performances on these benchmarks. The results of benchmarks evaluating general, coding, and math capacities are shown in Table 14. Table 15 and Table 16 show the performance on commonsense question answering and reading comprehension benchmarks respectively.
Model CW PE Data Method MMLU BBH LAMBADA HUMANEVAL MBPP GSM8K MATH
Llama-3-8B 8K - - - 64.80 63.97 75.78 34.75 48.28 50.04 10.20
32K ABF Long CPT 62.00 59.90 74.27 14.02 44.62 49.20 9.80
32K ABF Mix CPT 62.60 60.54 74.97 16.46 38.34 49.43 9.20
32K ABF Mix LongReD-C 64.40 62.78 75.35 35.37 44.42 52.84 10.20
32K ABF Mix LongReD-U 64.40 61.75 75.02 34.76 44.66 52.69 9.20
32K PI Long CPT 61.40 56.20 73.92 12.20 42.06 44.96 7.40
32K PI Mix CPT 60.50 57.17 74.99 16.46 38.12 49.13 9.20
32K PI Mix LongReD-C 63.30 60.83 75.22 34.76 42.02 51.55 10.00
128K ABF Long CPT 61.40 56.22 74.56 17.07 43.00 47.23 8.20
128K ABF Mix CPT 62.40 61.62 75.16 21.95 6.12 50.04 8.00
128K ABF Mix LongReD-C 63.60 60.25 74.11 35.37 44.68 52.92 7.60
128K ABF Mix LongReD-U 63.60 59.45 73.78 35.96 43.78 49.13 10.20
Mistral -7B-v0.3 32K - - - 62.30 58.49 75.28 24.39 42.00 45.56 8.40
128K ABF Long CPT 51.90 43.3 70.68 21.34 31.4 26.38 5.20
128K ABF Mix CPT 54.70 44.87 68.39 19.51 33.2 27.60 3.00
128K ABF Mix LongReD-C 58.90 51.58 73.86 20.12 33.2 33.36 6.40
128K ABF Mix LongReD-U 60.1 52.08 75.1 23.78 35.5 34.12 5.00
Table 14: Comparison of performances of short-text and long-text benchmarks of our methods with other baselines. CW denotes the context window length, PE denotes the scaling method of RoPE, RC denotes reading comprehension, and Common denotes commonsense question answering.
Model CW PE Data Method CommonsenseQA OpenBookQA PIQA SIQA ARC-C
Llama-3-8B 8K - - - 73.22 71 77.97 61 79.52
32K ABF Long CPT 71.50 66.40 74.1 59.42 77.73
32K ABF Mix CPT 72.65 69.60 77.09 62.74 77.30
32K ABF Mix LongReD-C 72.56 71.80 77.80 60.34 79.27
32K ABF Mix LongReD-U 72.89 71.80 77.91 61.46 79.01
32K PI Long CPT 71.25 62.80 72.03 55.58 76.88
32K PI Mix CPT 70.19 69.00 77.04 62.28 76.96
32K PI Mix LongReD-C 71.91 70.80 75.90 59.98 77.90
128K ABF Long CPT 65.19 64.00 77.15 59.83 75.68
128K ABF Mix CPT 72.65 70.00 77.53 61.62 77.73
128K ABF Mix LongReD-C 72.40 69.20 76.01 61.16 78.75
128K ABF Mix LongReD-U 72.40 69.80 77.04 61.00 77.99
Mistral -7B-v0.3 32K - - - 71.50 66.8 66.65 56.4 79.01
128K ABF Long CPT 66.99 34.8 56.8 38.89 67.15
128K ABF Mix CPT 66.99 34.8 56.8 38.89 67.15
128K ABF Mix LongReD-C 69.04 65.4 62.68 53.89 75.26
128K ABF Mix LongReD-C 70.84 65.6 63.87 55.48 78.5
Table 15: Comparison of performances of short-text and long-text benchmarks of our methods with other baselines. CW denotes the context window length, PE denotes the scaling method of RoPE, RC denotes reading comprehension, and Common denotes commonsense question answering.
Model CW PE Data Method SquadV2 quacQ TriviaQA BoolQ DROP
Llama-3-8B 8K - - - 72.06 35.65 75.42 82.6 53.18
32K ABF Long CPT 66.56 34.03 70.67 83.12 50.52
32K ABF Mix CPT 0.89 16.97 72.60 82.60 0.17
32K ABF Mix LongReD-C 74.18 34.39 74.24 82.29 49.10
32K ABF Mix LongReD-U 72.45 34.50 74.16 82.26 49.18
32K PI Long CPT 73.26 33.09 70.95 81.16 51.29
32K PI Mix CPT 7.53 21.43 72.55 81.44 0.21
32K PI Mix LongReD-C 71.96 34.73 74.87 82.94 49.47
128K ABF Long CPT 66.01 32.81 73.13 81.16 49.35
128K ABF Mix CPT 1.08 18.41 74.06 82.66 0.21
128K ABF Mix LongReD-C 70.67 33.97 75.19 82.51 49.63
128K ABF Mix LongReD-U 70.26 34.85 75.13 82.84 49.29
Mistral -7B-v0.3 32K - - - 72.47 31.84 76.1 83.43 45.46
128K ABF Long CPT 55.15 29.07 62.9 80.83 37.4
128K ABF Mix CPT 53.92 31.88 63.85 78.32 35.73
128K ABF Mix LongReD-C 70.45 30.16 68.16 80.61 41.9
128K ABF Mix LongReD-U 72.6 30.82 71.84 81.16 43.5
Table 16: Comparison of performances of short-text and long-text benchmarks of our methods with other baselines. CW denotes the context window length, PE denotes the scaling method of RoPE, RC denotes reading comprehension, and Common denotes commonsense question answering.

### v3

@@
 (2023); Dubey et al.
+(2024); Tang et al.
+(2025c); Wang et al.
+(2025); Chen et al.
 (2025).
@@
 (2023b); Dong et al.
-(2024).
+(2024a).
 To address the challenges of long context processing, various methods have been proposed to extend the context window of LLMs bloc97 (2023); Xiao et al.
@@
 However, as shown in Figure 1, these long context window extension techniques come at the cost of degraded performance on short-text tasks Xiong et al.
-(2024).
+(2024); Dong et al.
+(2024b).
 Yet, the root causes of this performance decline and potential mitigation strategies remain under-explored.
@@
 (2023); Dubey et al.
-(2024).
+(2024); Tang et al.
+(2025b).
 Given a Transformer decoder with $L$ layers and $N$ heads and an input sequence $\bm{x}=\{x_{1},\dots,x_{T}\}$ consisting with $T$ tokens, the output of the $l$-th layer can be denoted as $\mathbf{H}_{l}=\{\bm{h}_{l,1},\dots,\bm{h}_{l,T}\}$.
@@
 Thus, we only select $M(<L)$ layers as the distill layers $\mathbf{L}_{M}=\{l_{1},\dots,l_{M}\},\forall l_{i}\in\{0,\dots,L\}$ and distill the outputs of these layers.
-Specifically, we propose an attention-based layer selection method where we first compute the KL divergence of attention scores between the extended and original models and then select layers with the largest KL divergences for distillation.
-We also select the critical last layer that directly determines the output Men et al.
+Specifically, we propose an attention-based layer selection method where we first compute the KL divergence of attention scores between the extended and original models and then select layers with the largest KL divergences for distillation Men et al.
+(2024); Dong et al.
 (2025).
+We also select the critical last layer that directly determines the output.
 Short-Text Distillation Loss.
@@
 In addition to long-text training, our method incorporates both short-text distillation and short-to-long distillation.
-To evaluate the individual contributions of these two components, we employ half of the training data for long-text training and another half data for short-text distillation or short-to-long distillation.
+To evaluate the individual contributions of these two components, we employ half of the training data for long-text training and the other half for short-text distillation or short-to-long distillation.
 Moreover, we also evaluate the effect of the hyper-parameters $\alpha_{1}$ and $\alpha_{2}$ on both short and long text tasks.
@@
 Comparison with Continual Learning Methods.
-Beyond simple continual pre-training, we also compare LongReD with two continual learning methods that can effectively mitigate catastrophic forgetting issues: (1) Model Merging: Following previous work Hu et al.
-(2024b), we average the checkpoints of the original model and the extended model; (2) Parameter-Efficient Tuning: Since modifying RoPE directly affects attention computation without altering the knowledge stored in MLP parameters, we only tune the parameters in the attention blocks.
-We train Llama-3-8B on a 32K context length with these methods and evaluate their performances.
+Beyond simple continual pre-training, we also compare LongReD with two continual learning methods that can effectively mitigate catastrophic forgetting issues: (1) Model Merging: Following common model merging methods Hu et al.
+(2024b); Chen et al.
+(2024b), we average the checkpoints of the original model and the extended model; (2) Parameter-Efficient Tuning: Since modifying RoPE directly affects attention computation without altering the knowledge stored in MLP parameters, we only tune the parameters in the attention blocks Chen et al.
+(2024a).
+We train Llama-3-8B on a 32K context length with these methods and evaluate their performance.
 As shown in Table 7, though these methods can achieve better performances than naive continual pre-training, our method consistently performs better on most short-context capacities evaluations.
+This may be because these methods only address catastrophic forgetting.
+However, our approach not only addresses catastrophic forgetting through continual learning but also leverages knowledge distillation to alleviate distribution drift.
 Method General Code Math Common RC RULER
@@
 (2022); Han et al.
+(2023); Dong et al.
 (2023).
-In order to meet the growing demand for processing long texts, various methods have been proposed to extend the context windows of LLMs based on modification of RoPE Su et al.
+In order to meet the growing demand for processing long texts Wang et al.
+(2024); Tang et al.
+(2025a); Peng et al.
+(2025), various methods have been proposed to extend the context windows of LLMs based on modification of RoPE Su et al.
 (2024).
-Positional Interpolation down-scale positional indices to prevent out-of-distribution positions.
+Positional Interpolation down-scales positional indices to prevent out-of-distribution positions.
 In addition, NTK bloc97 (2023); emozilla (2023), ABF Xiong et al.
@@
 A more refined extension method that minimizes disruption to the model’s distribution could further enhance the effectiveness of our approach, a topic we leave for future investigation.
+Acknowledgement
+This work was partially supported by National Natural Science Foundation of China under Grant No.
+92470205 and 62222215, Beijing Municipal Science and Technology Project under Grant No.
+Z231100010323009, and Beijing Natural Science Foundation under Grant No.
+L233008.
+Xin Zhao is the corresponding author.
 Appendix A Skipped Positional Indices
@@
 For RULER, the sample is correct if the answer is in the output or each part of the answer (only for qa1 and qa2).
+We evaluate these models using the LLMBox Tang et al. (2024) library for most benchmarks, and OpenCompass Contributors (2023) for the math-related benchmarks.
 Appendix D Models for Evaluating Distribution Drift
@@
 Appendix E Analysis of Distillation Length
-As discussed in Section 5.3, increasing the distillation length will harm the long text modeling capacities.
-To explore the reason for the performance decline, we employ positional vectors Dong et al.
-(2024), a method to extract latent positional information from the hidden states.
-Specifically, given the hidden states from different samples, we average the hidden states at the same position to obtain vectors that representing positional information, denoted as positional vectors:
-$\bm{p}_{l,i}=\frac{1}{N}\sum_{s=1}^{N}\bm{h}_{l,i}^{s}$ (40)
-Then, we compute the cosine similarities of positional vectors with different places.
-All the positional vectors are calculated on samples from SlimPajama Soboleva et al.
-(2023) with a length of 32000 tokens.
-The similarity matrices of the positional vectors of models with distillation lengths of 1024, 2048, and 8192 as well as the models trained with only long texts are shown in Figure 5.
-We can observe that a large distillation length will cause a significant discontinuity in the implicit positional information inside and outside the original context window.
-Figure 5: Simialrity matrices of positional vectors inside and outside the context window.
-Appendix F Results Details
-In this section, we display the details of performances on these benchmarks.
-The results of benchmarks evaluating general, coding, and math capacities are shown in Table 14.
-Table 15 and Table 16 show the performance on commonsense question answering and reading comprehension benchmarks respectively.
 Model CW PE Data Method MMLU BBH LAMBADA HUMANEVAL MBPP GSM8K MATH
@@
 CW denotes the context window length, PE denotes the scaling method of RoPE, RC denotes reading comprehension, and Common denotes commonsense question answering.
+As discussed in Section 5.3, increasing the distillation length will harm the long text modeling capacities.
+To explore the reason for the performance decline, we employ positional vectors Dong et al.
+(2024a), a method to extract latent positional information from the hidden states.
+Specifically, given the hidden states from different samples, we average the hidden states at the same position to obtain vectors that representing positional information, denoted as positional vectors:
+$\bm{p}_{l,i}=\frac{1}{N}\sum_{s=1}^{N}\bm{h}_{l,i}^{s}$ (40)
+Then, we compute the cosine similarities of positional vectors with different places.
+All the positional vectors are calculated on samples from SlimPajama Soboleva et al.
+(2023) with a length of 32000 tokens.
+The similarity matrices of the positional vectors of models with distillation lengths of 1024, 2048, and 8192 as well as the models trained with only long texts are shown in Figure 5.
+We can observe that a large distillation length will cause a significant discontinuity in the implicit positional information inside and outside the original context window.
+Figure 5: Simialrity matrices of positional vectors inside and outside the context window.
+Appendix F Results Details
+In this section, we display the details of performances on these benchmarks.
+The results of benchmarks evaluating general, coding, and math capacities are shown in Table 14.
+Table 15 and Table 16 show the performance on commonsense question answering and reading comprehension benchmarks respectively.
 Model CW PE Data Method CommonsenseQA OpenBookQA PIQA SIQA ARC-C